OpenStudy (anonymous):

How do I solve for the limit of x->0, xsin(1/16pix)

OpenStudy (anonymous):

substitue 0 directly into the limit

OpenStudy (anonymous):

change \[\lim_{x \rightarrow 0} x.\sin(pix/16)\] into quotient in order to use L'Hopital's Rule \[\lim_{x \rightarrow 0} \sin(pix/16)/1/x\] now differentiate numerator and denominator separately , you get: \[\lim_{x \rightarrow 0} (\pi/16)\cos(pix/16)/1/1\] now you can apply the limit, as x approaches 0, y = \[\pi/16\]